Analysis of a 10.15g sample of a compound known to contain only phosphorous and oxygen is found to contain 5.717g of oxygen. What is the simplest formula of this compound?
1
Expert's answer
2020-03-16T12:17:23-0400

The number of moles of oxygen in the following compound is:  

"n(O)=m(O)\/M(O)"

"n(O) = 5,717 g\/(16g\/mol) = 0,357 mol"

If the compound contains only phosphorous and oxygen then the rest of the mass is the mass of phosphorous:

"m(P)=10.15g-5.717g=4,433g"

"n(P)=m(P)\/M(P)"

"n(P)=4,433g\/(31g\/mol)=0,143 mol"

The formula of a compound is P0,143O0,357

Let us divide 0,357 by 0,143

"0,357\/0,143=2.5"

Now the formula is PO2,5

We can get the closest integers of atoms by multiplying by 2

The final formula is P2O5